# LieRE: Lie Rotational Positional Encodings

### v3

@@ Abstract
Transformer architectures rely on position encodings to capture token dependencies. Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations. However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity. We address these challenges with Lie Relative Encodings (LieRE), which replaces RoPE’s block-2D rotation matrix with a learned, dense, high-dimensional rotation matrix of variable density. Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 2% relative improvement over state-of-the-art baselines on 2D tasks and 1.5% on 3D tasks, while demonstrating superior generalization to higher resolutions. Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100, and we release our code to facilitate further research.
@@ 1 Introduction
The attention mechanism, particularly within transformer architectures, has revolutionized machine learning across diverse domains. However, attention is inherently permutation—invariant—it cannot leverage the sequential order of its inputs as information. This fundamental limitation necessitates additional mechanisms to encode positional information, enabling models to capture sequential and spatial dependencies crucial for tasks ranging from natural language processing to image understanding (Vaswani et al., 2017).
@@
Figure 1: Overview of LieRE, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$. $d$ is the head dimension. The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
@@
To assess the impact of LieRE and other positional encodings on ViT performance, we evaluate several encoding schemes across diverse tasks, including 2D and 3D image classification. Additionally, we investigate a fundamental spatial reasoning task where models must identify arrow directions. Despite their sophistication, contemporary multimodal LLMs struggle with this seemingly simple task. Our experiments reveal that successful completion of this task specifically requires relative position encodings, highlighting their crucial role in spatial understanding.
@@ 2.1 Position Encodings
We organize our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
Absolute position in this context refers to a position with respect to a consistent reference, usually the start of a tex or top left corner of an image. Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the input. Methods such as sinusoidal and learned absolute encodings directly add vectors to the input token embedding (Vaswani et al., 2017; Devlin et al., 2019; Dosovitskiy et al., 2020).
Relative position encodings in contrast encode the relative positions of two tokens. One strategy is to learn an set of embeddings for position deltas which can be incorporated into the attention mechanism (Shaw et al., 2018b; Liu et al., 2021, 2022). However, this incurs quadratic computational cost in terms of the number of tokens as a separate embedding is required for every pair of tokens. Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product. The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism (Su et al., 2024). RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models (Touvron et al., 2023; Chowdhery et al., 2023; Jiang et al., 2024). However, RoPE can perform poorly on inference for larger context sizes than the model was trained on. This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
@@ 3.1 Lie Groups in the Context of Attention
Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks 4. These form a special Lie group that is commutative, allowing us to strengthen the statement in (1) to
@@
Figure 3: Pseudocode for the Lie Rotary Embedding (LieRE) attention mechanism. Unlike RoPE, LieRE allows for dense rotation matrices.
@@ 4 Method
LieRE is a simple modification to the attention mechanism that is presented in algorithm 3. Recall that we assume that positions are $n$-dimensional vectors, a matrix $A$ is skew-symmetric if $A^{T}=-A$, and that the matrix exponential of a skew-symmetric matrix, call it $A$, is always a high dimensional rotation matrix.
@@
Figure 4: Pseudocode for the Rotary Position Embedding (RoPE) attention mechanism.
@@
We include the pseudocode for the LieRE attention in algorithm 3 in addition to the standard RoPE attention (algorithm 4). In practice, we compute the rotation matrices at the start of the forward pass. By default, the skew symmetric bases are learned separately for every layer and attention head except in the experimental section focused on sharing parameters across heads and layers.
5 Experiments
In order to isolate the effect of changing the position encoding, we use a plain standard transformer backbone and training regime [^1] https://github.com/kentaroy47/vision-transformers-cifar10 with comparable baseline performance on CIFAR100 and ImageNet with (Lee et al., 2022). We implement relative position encoding types including LieRE, Rope-Mixed, Visionllama, and absolute position encodings. We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L (Dosovitskiy et al., 2020). All experiments use RandAugment (Cubuk et al., 2020). We avoid using pre-trained weights in order to maximize the comparability of results between methods. In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B). We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively. Notably, a tile size of 2 corresponds to RoPE-Mixed.
5.1 2D Classification
For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task (Krizhevsky et al., 2009; Deng et al., 2009). We further analyse different backbone models sizes and data ablation effects. We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024). For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (table 1).
Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results. All models use 85.2M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT architecture, ∗∗ equivalent to Vivit (spatio-temporal) architecture.
@@ 5 Experiments
We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020). Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in table 5. Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
@@ 5.2 Synthetic Spatial Classification
Figure 7: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down.
@@
We train the models on 800,000 examples and observe that they generally converge after the first 400,000. The only exception to this is absolute position encodings, where we have variants trained on 800,000 and
@@
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-Mixed and absolute encoding similar to the previous section (Arnab et al., 2021; Heo et al., 2024). For the 3D experiments, we examine video classification performance in the UCF101 dataset (Soomro et al., 2012). Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022). The full set of hyperparameters may be found in appendix B. We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
5.4 LieRE parameter
LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is. To answer this question, we vary the density of the learned basis and the learnability across heads and layers.
We vary capacity by enforcing a block diagonal structure on the skew-symmetric basis. Varying the block size allows us to approach LieRE64. Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed. As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024). In figure 8 we evaluate accuracy versus block dimension (Appendix Table 8).
@@ 5.6 Multi-resolution Classification
In this section we compare the ability of methods to generalize to image resolutions not seen during training. We evaluate two training recipes inspired by (Heo et al., 2024). The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs. The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs. The full details can be found in appendix B. We evaluate the accuracy on the ImageNet validation set with varying inference resolutions. Specifically, we scale the input images to resolutions of $196\times 196$, $256\times 256$, $320\times 320$, $384\times 384$, and $448\times 448$ pixels per dimension, and present the resulting accuracies in figure 9. For position assignment, we adopt a sequential approach where token positions are scaled proportionally to the image dimensions. For example, doubling the length of an image in each dimension doubles the range of positional indices. This method outperforms rescaling positions to a fixed range, as demonstrated by superior results for both RoPE-Mixed and LieRE across the evaluated training recipes.
@@ 7 Limitations
8 Broader Impacts
LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated improvements in 2D image and video classification tasks. We are particularly excited about how LieRE can expand the applicability transformers to $n$ dimensional inputs. This may lay down the tracks to apply that same relative position encoding within and across modalities and settings.
Appendix A Appendix
Appendix B Experiment Hyperparameters

### v4

@@ Abstract
Transformer architectures depend on explicit position encodings to capture token positional information. Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations. However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity. We address these challenges with Lie Relative Encodings (LieRE), which generalizes RoPE to high-dimensional rotation matrices by leveraging their Lie group structure. Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 1.5% improvement over state-of-the-art baselines on 2D tasks and 1% on 3D tasks, while demonstrating superior generalization to higher resolutions. Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100. Our code is available at https://github.com/StanfordMIMI/LieRE.
@@ 1 Introduction
The attention mechanism, particularly within transformer architectures, has revolutionized machine learning across diverse domains. However, attention is inherently permutation invariant—it cannot leverage the sequential order of its inputs as information directly. This fundamental limitation necessitates additional mechanisms to encode positional information, enabling models to capture sequential and spatial dependencies crucial for tasks ranging from natural language to image understanding (Vaswani et al., 2017).
@@
Figure 1: Overview of how LieRE encodes spatial information, where $N$ denotes the number of input dimensions, i.e. $N=3$ for the 3D image. $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x_{i}\;y_{i}\;z_{i}])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the $i$th patch in the flattened input, where $j$ corresponds to a different patch. $d$ is the head dimension.
@@
To assess the impact of LieRE and other positional encodings on ViT performance, we evaluate several encoding schemes across diverse tasks, including 2D and 3D image classification. Additionally, we investigate a fundamental spatial reasoning task where models must identify where an arrow points. Despite their sophistication, contemporary multimodal LLMs struggle with this seemingly simple task. Our experiments reveal that successful completion of this task specifically requires relative position encodings, highlighting their crucial role in spatial understanding.
@@ 2.1 Position Encodings
We split the review of positional encodings into: a) absolute, b) relative, and c) contextual.
Absolute position in this context refers to a position with respect to a consistent reference, usually the start of a text or top left corner of an image. Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the input. Methods such as sinusoidal and learned absolute encodings directly add vectors to the input token embedding (Vaswani et al., 2017; Devlin et al., 2019; Dosovitskiy et al., 2020).
Relative position encodings, in contrast, encode the relative positions of two tokens. One strategy is to learn an set of embeddings for position deltas which can be incorporated into the attention mechanism (Shaw et al., 2018b; Liu et al., 2021, 2022). However, this incurs quadratic computational cost in terms of the number of tokens as a separate embedding is required for every pair of tokens. Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product. The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism (Su et al., 2024). RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models (Touvron et al., 2023; Chowdhery et al., 2023; Jiang et al., 2024). However, RoPE can perform poorly on inference for larger context sizes than the model was trained on. This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
@@ 2.2 Extensions of RoPE
2.3 Efficient Scaling beyond Sequence Data
There is also an extensive line of work improving the performance of transformers for modalities with dimensionality greater than one. Axial Attention (Ho et al., 2019) reduces computational complexity by applying attention along specific axes (e.g., rows and columns in images), enabling transformers to scale efficiently with high-dimensional data. Perceiver utilizes latent tokens to compress high-dimensional inputs into a smaller set, improving scalability as input size and dimensionality grow. These methods address the inefficiencies of traditional transformers when applied to high-dimensional data (Jaegle et al., 2021). Additionally, techniques like Swin and Vmamba optimize compute for visual data through structuring how information flows through the network (Liu et al., 2021, 2024). Swin Transformer introduces a hierarchical approach with shifted windows, limiting attention to local regions to reduce complexity while capturing global context. Vmamba, on the other hand, proposes a visual state space model that represents images as a collection of spatial states, allowing attention to be applied efficiently across large-scale visual inputs by exploiting spatial locality and reducing redundant computation.
2.4 Equivariant Networks
A related branch of work encoding problem structure focuses on equivariance. We say that a model $T$ is equivariant with respect to if $T(f(x))=g(T(x))g$ (Tai et al., 2019). Where with relative position encoding we often want to be able to encode translation invariance, equivariance provides a more general framework. Equivariance has been applied to improve performance on problems with a wide array of structures, ranging from rotation-invariance (Esteves et al., 2017, 2018; Worrall et al., 2017), 3D reference frame-invariance (Liao & Smidt, 2022; Fuchs et al., 2020) and many others. The subset of these works that focus on generating equivariant token embeddings for transformers can be combined directly with LieRE or another rotation-based position encoding.
2.5 Lie Groups in Machine Learning
Lie groups have also had extensive use in machine learning. The range of works is diverse, ranging from algebraic signal processing (Kumar et al., 2024), automated discovery of symmetries (Forestano et al., 2023) to state estimation (Falorsi et al., 2019). Furthermore (Gallier & Quaintance, 2020) provides a friendly introduction to differential geometry and lie groups that may be of interest to the reader.
@@ 3.1 Lie Groups in the Context of Attention
Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks 3(b). These form a special Lie group that is commutative, allowing us to strengthen the statement in (1) to
@@
(a) Lie Rotary Embedding (LieRE) attention mechanism.
(b) Rotary Position Embedding (RoPE) attention mechanism.
Figure 3: Comparison of the LieRE and RoPE-Mixed attention mechanisms.
@@ 4 Method
LieRE is a simple modification to the attention mechanism that is presented in Algorithm 3(a). Recall that we assume that positions are $n$-dimensional vectors, a matrix $A$ is skew-symmetric if $A^{T}=-A$, and that the matrix exponential of a skew-symmetric matrix, call it $A$, is always a high dimensional rotation matrix.
@@
We include the pseudocode for the LieRE attention in Algorithm 3(a) in addition to the standard RoPE attention (Algorithm 3(b)). In practice, we compute the rotation matrices at the start of the forward pass. By default, the skew symmetric bases are learned separately for every layer and attention head except in the experimental section focused on sharing parameters across heads and layers.
Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results. All models use 85.2M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Flanders et al., 2020) equivalent to DeiT architecture, ∗∗ equivalent to Vivit (spatio-temporal) architecture.
@@
Table 2: P-values comparing RoPE-Mixed and LieRE8 across datasets.
Dataset P-value (RoPE-Mixed vs. LieRE8)
CIFAR-100 $1.3\times 10^{-5}$
ImageNet-1k $6.3\times 10^{-3}$
UCF101 $7.1\times 10^{-4}$
$384\times 384$ (Resolution Invariance) $4.0\times 10^{-4}$
5 Experiments
In order to isolate the effect of changing the position encoding, we use a plain standard transformer backbone and training regime [^1] https://github.com/kentaroy47/vision-transformers-cifar10 with comparable baseline performance on CIFAR100 and ImageNet with (Lee et al., 2022). We implement relative position encoding types including LieRE, RoPE-Mixed, Visionllama, and absolute position encodings. We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L (Dosovitskiy et al., 2020). All experiments use RandAugment (Cubuk et al., 2020). We avoid using pre-trained weights in order to maximize the comparability of results between methods. In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B.1). We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively. Notably, a tile size of 2 corresponds to RoPE-Mixed.
5.1 2D Classification
For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task (Krizhevsky et al., 2009; Deng et al., 2009). We further analyse different backbone models sizes and data ablation effects. We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024). For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (Table 1, Table 2).
We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020). Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in Table 6, Figure 8. Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
@@ 5.2 Synthetic Spatial Classification
Figure 5: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down. This task requires understanding spatial relationships.
@@
We train the models on 800,000 examples and observe that they generally converge after the first 400,000 examples. The only exception to this is absolute position encodings, where all variants were trained on 800,000 examples with no sign of increases in performance.
@@
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-Mixed and absolute encoding similar to the previous section (Arnab et al., 2021; Heo et al., 2024). For the 3D experiments, we examine video classification performance in the UCF101 dataset (Soomro et al., 2012). Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022). The full set of hyperparameters may be found in appendix B.1. We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
5.4 LieRE Capacity
LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is. To answer this question, we vary the density of the learned basis and the learnability between heads and layers.
We vary capacity by enforcing a block-diagonal structure on the skew-symmetric basis. Varying the block size allows us to approach a dense matrix (LieRE64 for our models). Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed. As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024). We found that RoPE-Mixed is more sensitive to the initialization method than LieRE as seen in appendix table 7. In figure 6 we evaluate accuracy versus block dimension (Appendix Table 10).
@@ 5.6 Multi-resolution Classification
In this section we compare the ability of methods to generalize to image resolutions not seen during training. We evaluate two training recipes inspired by (Heo et al., 2024). The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs. The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs. The full details can be found in appendix B.1. We evaluate the accuracy on the ImageNet validation set with varying inference resolutions. Specifically, we scale the input images to resolutions of $196\times 196$, $256\times 256$, $320\times 320$, $384\times 384$, and $448\times 448$ pixels per dimension, and present the resulting accuracies in figure 7. For position assignment, we adopt a sequential approach where token positions are scaled proportionally to the image dimensions. For example, doubling the length of an image in each dimension doubles the range of positional indices. This method outperforms rescaling positions to a fixed range, as demonstrated by superior results for both RoPE-Mixed and LieRE across the evaluated training recipes.
@@ 7 Limitations
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
We would like to thank Maksim Maydanskiy for helping us understand Lie groups and Lie group representations. We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper. Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550). This project was supported by Google Cloud and Azure credits.
Appendix A Equivalence of RoPE and LieRE in one Dimension
Though focused on higher dimensional inputs LieRE remains compatible with 1D tasks. It turns out that in the 1D setting, LieRE has identical representational capacity to RoPE. This is not the case for higher dimensional inputs
Recall that in the 1D setting positions are scalars. The LieRE rotation is $R=\exp(tA)$ for a some learnable skew-symmetric matrix $A$. Recall that skew-symmetric matrices can be written in the form $S^{T}\Lambda S$ where and is orthogonal $\Lambda$ has the structure denoted below.
$\Lambda=\begin{pmatrix}0&\lambda_{0}&&&&\\ -\lambda_{0}&0&&&&\\ &&0&\lambda_{1}&&\\ &&-\lambda_{1}&0&&\\ &&&&\ddots&\\ \end{pmatrix}$
We can then use an identity of the matrix exponential to break down the LieRE rotation matrix.
$R=\exp(tS^{T}\Lambda S)=S^{T}\exp(t\Lambda)S.$
For two tokens in positions $t,t^{\prime}$ we denote the embeddings for a specific attention head as $x_{t},x_{t^{\prime}}$. If $K,Q$ denote the corresponding key and query linear transformation matrices we can write the attention inner product with LieRE explicitly.
$x_{t}KR_{t}^{T}R_{t^{\prime}}Qx_{t^{\prime}}$ $=x_{t}KS^{T}\exp(t\Lambda)S^{T}S\exp(t^{\prime}\Lambda)SQx_{t^{\prime}}$
$=x_{t}KS^{T}\exp(t\Lambda)^{T}\exp(t^{\prime}\Lambda)SQx_{t^{\prime}}$
$=x_{t}KS^{T}\exp(t\Lambda)^{T}\exp(t^{\prime}\Lambda)SQx_{t^{\prime}}$
We let $K^{\prime}=KS^{T}$ and $Q^{\prime}=SQ$, since these matrices are all learnable we can fold the $S$ matrix into parameters of the key and query linear layers for the given head, allowing us to simplify the above expression.
$x_{t}K^{\prime}\exp(t\Lambda)^{T}\exp(t^{\prime}\Lambda)Q^{\prime}x_{t^{\prime}}$
Now we use the fact that each block is skew symmetric. In the case of two dimensions,
$\exp\left(\begin{pmatrix}0&\lambda\\ -\lambda&0\end{pmatrix}\right)=\begin{pmatrix}\cos(\lambda)&\sin(\lambda)\\ -\sin(\lambda)&\cos(\lambda)\end{pmatrix}$
If we let $R_{\lambda}$ denote a block diagonal rotation matrices with 2D rotations of angles $\lambda_{0},\ldots,\lambda_{n}$, we can rewrite the above expression in a more familiar form.
$x_{t}K^{\prime}R_{t}^{T}R_{t^{\prime}}Q^{\prime}x_{t^{\prime}}$
This is exactly the formulation of the original RoPE position encoding. This also makes more clear how LieRE is different from RoPE-Mixed in the high dimensional setting. The above proof depends on the fact that we can decompose every rotation into a matrix of the form $S^{T}\Lambda S$ with $S$ not dependent on the position, allowing us to fold the orthogonal $S$ matrices into the key and query matrices. This decomposition with constant $S$ is guaranteed because the inputs to the matrix exponential differ by only a scalar factor. This is no longer true once we switch to more than a one-dimensional basis of skew symmetric matrices.
Appendix B Experimental Details
B.1 Experimental Hyperparameters